how do i solve this? : 4/9 of 81? thankies ^.^
Answer by jim_thompson5910(35256)

4/9 of 81 = (4/9)*(81)

4/9 of 81 = (4/9)*(81/1)

4/9 of 81 = (4*81)/(9*1)

4/9 of 81 = (4*9*9)/(9*1)

4/9 of 81 = (4*9)/(1)

4/9 of 81 = (36)/(1)

4/9 of 81 = 36
--------------------------------------------------------------------------------------------------------------
If you need more help, email me at jim_thompson5910@hotmail.com

Also, please consider visiting my website: http://www.freewebs.com/jimthompson5910/home.html and making a donation. Thank you

Jim
--------------------------------------------------------------------------------------------------------------